OpenStudy (linda):

express in radical form : (sin pi over 4) (tan pi over 6) - (tan pi over 3) (cos pi over 2) plz help

OpenStudy (watchmath):

\(\frac{\sqrt{2}}{{2}}\cdot \frac{1}{\sqrt{3}}-\sqrt{3}\cdot 0\) You continue from there

OpenStudy (anonymous):

Remember to use parenthesis! I was confused about whether or not you meant sin(pi)/4 or sin(pi/4) for example.

OpenStudy (linda):

oh ok sorry! so when i multiply square root of 2 over 2 times 1 over square root of 3 i get square root of 2 over s square root of 3...

OpenStudy (linda):

but i know you cant have a radical on the denominator so would i multiply the numerator by 2 square root of 3?

OpenStudy (anonymous):

you have to multiply fractions by one always when you are simplifying, otherwise you are changing the fraction's value! So multiply it by something/something. What would the something be to get rid of the square root of 2? Remember that: \[\sqrt{x}*\sqrt{x} = x\] (if x is positive)

OpenStudy (linda):

it would just be square root of 2 right?